Marshall Dillon is riding at 30 mph after the robber of the Dodge City bank, who has a head startof 15 minutes, but whose horse can only make 25 mph on a good day. How long does it take forDillon to catch up with the bad guy, and how far from Dodge Cityare they when this happens?(Assume the road is straight, for simplicity.
Answer #1

The robber had a head start by

25 mph * (15/60) = 6.25 miles

let d be the distance covered by dillon then

d - 6.25 = 25 * t (robber's equation)

d = 30t ( dillon's equation)

set both equations equal

25t + 6.25 = 30t

6.25 = 5t

t = 1.25 hours

--------------------------------------------------------------

To find the distance put value of time in either equations above

d = 30t

d = 30 * 1.25

d = 37.5 miles